# Convolutional and Deep Learning based techniques for Time Series Ordinal Classification

### v2

@@ Appendix A Data description
First of all, we made a dataset selection from two time series repositories, the UEA/UCR TSC archive [^1] https://timeseriesclassification.com, and the Monash/UEA/UCR TSER archive [^2] http://tseregression.org/. For the TSC repository, we analysed the output variable of each problem to select those with an ordinal nature. We have identified seven univariate, and two multivariate datasets. Moreover, we have chosen 13 additional datasets from the TSER archive. Given that it is a regression repository, we have discretised the original continuous output variable, grouping it into five uniformly distributed bins. Even though there are more regression datasets, we could only select these 13 as it is difficult to achieve a reasonable class distribution for the rest (the output values are too close to obtain more than three representative classes). Furthermore, the datasets with unequal length time series and missing values are also discarded. Other five TSOC problems built from historical price data from five of the most important companies in the stock market are considered: Apple (AAPL), Amazon (AMZN), Google (GOOG), META and Microsoft (MSFT), and other two were built from data collected from the National Data Buoy Center (NDBC).
Data for the five TSOC problems built from historical price data has been collected from Yahoo Finance [^3] https://es.finance.yahoo.com/ website, extracting weekly price data from each stock earliest available date to March 2023. The time series of each stock data are formed by its price returns over 53 weeks (around one year) before a varying date $t$. The output ordinal label corresponds to the price return in $t$. This value is discretised according to a set of predefined symmetrical thresholds $(-\infty,-0.05,-0.02,0.02,0.05,\infty)$. Finally, two more TSOC problems were built from data collected from the National Data Buoy Center (NDBC) [^4] https://www.ndbc.noaa.gov/. Input multivariate time series are seven climatological variables collected from a reanalysis model, as detailed in [gomez2022simultaneous]. Some of these variables are the air temperature, the pressure, or the relative humidity. Each time series comprises four values per day during four weeks, i.e. the length of the time series is $112$. The output label indicates the energy fluctuation level during the data collection period, with 00 being the minimum and $3$ the highest level. The second problem also comprises multivariate time series. Same seven climatological variables are considered for this problem. This time, the length of the time series is $28$, as collects values during only one week (four values per day). The purpose is to estimate the wave height level during that period of time, from 00 (the lowest height) to $3$ (the highest height). Note that the outputs of both datasets have been discretised in bins with identical widths.

### v3

@@ Appendix A Data description
First of all, we made a dataset selection from two time series repositories, the UCR TSC archive [^1] https://timeseriesclassification.com, and the Monash TSER archive [^2] http://tseregression.org/. For the TSC repository, we analysed the output variable of each problem to select those with an ordinal nature. We have identified seven univariate, and two multivariate datasets. Moreover, we have chosen 13 additional datasets from the TSER archive. Given that it is a regression repository, we have discretised the original continuous output variable, grouping it into five uniformly distributed bins. Even though there are more regression datasets, we could only select these 13 as it is difficult to achieve a reasonable class distribution for the rest (the output values are too close to obtain more than three representative classes). Furthermore, the datasets with unequal length time series and missing values are also discarded. Other five TSOC problems built from historical price data from five of the most important companies in the stock market are considered: Apple (AAPL), Amazon (AMZN), Google (GOOG), META and Microsoft (MSFT), and other two were built from data collected from the National Data Buoy Center (NDBC).
Data for the five TSOC problems built from historical price data has been collected from Yahoo Finance [^3] https://es.finance.yahoo.com/ website, extracting weekly price data from each stock earliest available date to March 2023. The time series of each stock data are formed by its price returns over 53 weeks (around one year) before a varying date $t$. The output ordinal label corresponds to the price return in $t$. This value is discretised according to a set of predefined symmetrical thresholds $(-\infty,-0.05,-0.02,0.02,0.05,+\infty)$. Finally, two more TSOC problems were built from data collected from the National Data Buoy Center (NDBC) [^4] https://www.ndbc.noaa.gov/. Input multivariate time series are seven climatological variables collected from a reanalysis model, as detailed in [gomez2022simultaneous]. Some of these variables are the air temperature, the pressure, or the relative humidity. Each time series comprises four values per day during four weeks, i.e. the length of the time series is $112$. The output label indicates the energy fluctuation level during the data collection period, with 00 being the minimum and $3$ the highest level. The second problem also comprises multivariate time series. Same seven climatological variables are considered for this problem. This time, the length of the time series is $28$, as collects values during only one week (four values per day). The purpose is to estimate the wave height level during that period of time, from 00 (the lowest height) to $3$ (the highest height). Note that the outputs of both datasets have been discretised in bins with identical widths.